OpenStudy (geekfromthefutur):

A sequence has its first term equal to 8 and each term of the sequence is obtained by adding 6 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 6 and f(n) = f(n - 1) + 8; n > 1 f(1) = 8 and f(n) = f(n - 1) + 6; n > 1 f(1) = 8 and f(n) = f(n - 1) + 6n; n > 1 f(1) = 6 and f(n) = f(n - 1) + 8n; n > 1

OpenStudy (phi):

A sequence has its first term equal to 8 can you rule out any of the 4 choices ?

OpenStudy (geekfromthefutur):

im thinking the third one

OpenStudy (phi):

Hopefully it is clear that A sequence has its first term equal to 8 means f(1)=8 so the first and last choices are definitely out.

OpenStudy (phi):

each term of the sequence is obtained by adding 6 to the previous term. we can list a few terms, to see what the sequence looks like. It starts at 8 8 and then add 6 to get the next number: 8+6= 14 then add 6 to that to get the third number: 14+6= 20 so 8,14,20,26, and so on.

OpenStudy (phi):

Let's try your choice: f(1) = 8 and f(n) = f(n - 1) + 6n; n > 1 when n=1, we get f(1)=8 (which is good) when n=2 we do f(2)= f(2-1) + 6*2 (replace n with 2 everywhere you see n) simplify to get: f(2) = f(1) + 12 f(1) is 8 (from the first step. so: f(2)= 8+12= 20 according to the rule, f(2)= 20 but the sequence goes 8,14,20,26... f(2) is 14. This rule is not the one we want.

OpenStudy (geekfromthefutur):

oh ok so it would be second one

OpenStudy (phi):

you can even check the 2nd choice: f(1) = 8 and f(n) = f(n - 1) + 6; n > 1 f(2)= f(2-1) + 6 f(2) = f(1) + 6 f(2) = 8+6 = 14 which works. This stuff is not too hard if you practice it.

OpenStudy (geekfromthefutur):

yea im not to great with math i try to study a lot but im not to good

OpenStudy (geekfromthefutur):

but is it ok if u could help me with one more question if possible?